OpenStudy (anonymous):

solve for x by factoring.(2x+1)(x+4)=0. show working out im struggling

OpenStudy (anonymous):

its 2x-1 my bad

OpenStudy (lgbasallote):

equate the factors to zero so... 2x -1 = 0; x+ 4 = 0 in the first factor... 2x = 1 so x = 1/2 in the second factor x = -4

OpenStudy (accessdenied):

It is fundamentally the same process as the previous problem you'd approach each problem in this form the same way. - factor completely (its already factored completely) - set each factor = 0